. Find the sonic velocity for the following fluids :

(i) Crude oil of specific gravity 0.8 and bulk modulus 1.5 GN/m2 ;

(ii) Mercury having a bulk modulus of 27 GN/m2.


1
Expert's answer
2021-09-21T06:34:04-0400

The speed of sound, according to the Newton–Laplace equation, is


"c_O=\sqrt{\frac{K_{sO}}{\rho_O}}=\sqrt{\frac{1.5·10^9}{0.8·1000}}=1369\text{ m/s}."

"c_M=\sqrt{\frac{K_{sM}}{\rho_M}}=\sqrt{\frac{27·10^9}{13600}}=1409\text{ m/s}."